# Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models

### v5

@@ C.2 Computing Value Correlations
The structure of a value system is derived from correlations between different values. In psychology research, researchers measure the value hierarchies of the participants to gather data, then use correlation derived from the data to evaluate interrelationships among these values. High positive correlations indicate values that are often endorsed together, while negative correlations show contrasting values. From these correlations, researchers can map and cluster values, revealing underlying value structures and hidden value factors (§ 3, Value Measurement and System Construction).
@@ Results.
Fig. 5 illustrates the correlation heatmap of Schwartz’s values based on the co-occurrence frequency of the lexicons. The values are ordered along the x and y axes according to Schwartz’s circular structure. Except for the two values of Self-Enhancement, correlations between values generally contradict the theoretical structure of Schwartz’s values. The results suggest that the co-occurrence frequency of lexicons in responses to direct prompts does not necessarily reflect the true value correlation. Value measurements using GPV are more theoretically grounded and empirically validated Ye et al. (2025b).
@@ C.3 LLM Subjects
Please note that value-anchoring prompts based on Schwartz’s values do not bias the structure of the value system. Theoretically, favoring a certain value (e.g., power) does not bias the structure of the value system (e.g., how self-direction is correlated with power). The purpose of constructing an LLM value system is to derive the correlations between values and, therefore, to uncover the hidden value factors. Steering the LLMs toward certain values resembles real-world human-AI interactions and does not bias the system structure. Additionally, building upon established theories is a common practice in psychology and can lead to novel structures (Cieciuch et al., 2014; Tevrüz et al., 2015). Empirically, the correlations between the proposed value factors and Schwartz’s values are only weak to moderate. The correlation between our values and Schwartz’s ten values is presented in Table 10; the aggregated correlation between our values and Schwartz’s four dimensions is presented in Table 11.

### v6

@@ C.2 Computing Value Correlations
The structure of a value system is derived from correlations between different values. In psychology research, researchers measure the value hierarchies of the participants to gather data, then use correlation derived from the data to evaluate interrelationships among these values. High positive correlations indicate values that are often endorsed together, while negative correlations show contrasting values. From these correlations, researchers can map and cluster values, revealing underlying value structures and hidden value factors (§˜3, Value Measurement and System Construction).
@@ Results.
Fig.˜9 illustrates the correlation heatmap of Schwartz’s values based on the co-occurrence frequency of the lexicons. The values are ordered along the x and y axes according to Schwartz’s circular structure. Except for the two values of Self-Enhancement, correlations between values generally contradict the theoretical structure of Schwartz’s values. The results suggest that the co-occurrence frequency of lexicons in responses to direct prompts does not necessarily reflect the true value correlation. Value measurements using GPV are more theoretically grounded and empirically validated Ye et al. (2025b).
@@ C.3 LLM Subjects
Please note that value-anchoring prompts based on Schwartz’s values do not bias the structure of the value system. Theoretically, favoring a certain value (e.g., power) does not bias the structure of the value system (e.g., how self-direction is correlated with power). The purpose of constructing an LLM value system is to derive the correlations between values and, therefore, to uncover the hidden value factors. Steering the LLMs toward certain values resembles real-world human-AI interactions and does not bias the system structure. Additionally, building upon established theories is a common practice in psychology and can lead to novel structures (Cieciuch et al., 2014; Tevrüz et al., 2015). Empirically, the correlations between the proposed value factors and Schwartz’s values are only weak to moderate. The correlation between our values and Schwartz’s ten values is presented in Table˜10; the aggregated correlation between our values and Schwartz’s four dimensions is presented in Table˜11.